OpenStudy (anonymous):

If 4x^2-y=100 and dy/dt =-6, find dx/dt when x=1.

OpenStudy (anonymous):

Differentiate.

OpenStudy (anonymous):

Can you give me a step-by-step?

OpenStudy (ranga):

First find dy/dx and put x = 1 and you will get a numerical value. dy/dx = dy/dt x dt/dx = (dy/dt) / (dx/dt) So dx/dt = (dy/dt) / (dy/dx) --- (1) When x = 1, dy/dt = -6 and you know dyx/dx at x = 1 from step 1. Substitute in (1) and compute dx/dt